• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name Zhihong Ye

    Log entry time 09:50:47 on March 08, 2012

    Entry number 364558

    keyword=G2p Status Update: Smooth

    1, Beamline
      MCC were able to deliver beam to the beam down after correcting the 5% 
    field offset of the second chicane before the target magnet. They spent a 
    while to figure that the reason they could not see the beam on the viewer 
    before beam dump was because the viewer itself went bad. 
      1H05 Harp (the one between second chicane and target magnet) is still  
    not functioning properly because of the large noise. We barely can see 
    the signal peaks from its wires so we still be cable to calibrate BPM and 
    slow raster.
      Alex,Pengjia and Toby were working hard during the owl shift to 
    calibrate BPM and check slow raster. We are waiting for their detail 
    2, We will take Optics data through rest of shifts before we move the 
    chicane again on Friday (Maybe), although we have not gotten data yet 
    since MCC has problem with one of their magnets and could not send the 
       Before we start taking optics data, Josh and James want to double 
    check the offsets of targets and do adjustment in target control system.

    A copy of this log entry has been emailed to: jpchen, slifer,camsonne,rom,doug